The perimeter of a certain rectangle is 80 meters, the length of the rectangle is 10 meters more than the width. Find the length and the width.
Answer by Boreal(15235)

Perimeter=2L+2W
Half perimeter is L+W=40
L=W+10
Therefore, W+10+W=40
2W=30
W=15 meters
L=25 meters